OpenStudy (anonymous):

two metal cubes with sides of 4 cm are melted down and casted into a spherical ball. Find the radius of the sphere so formed. formula for volume of cube: l x w x h formula for volume of sphere: 4/3(pi)r^2

OpenStudy (anonymous):

@nelsonjedi could you help me on this?

OpenStudy (anonymous):

What is the volume of the cube?

OpenStudy (anonymous):

64 cm^3

OpenStudy (anonymous):

So 4/3(pi)r^2 = 64

OpenStudy (anonymous):

lol. i was just about to say that. yeah.

OpenStudy (anonymous):

So we need to isolate r^2...How would we do that?

OpenStudy (anonymous):

divide by 4/3 and then divide by pi?

OpenStudy (anonymous):

Yes divide by 4/3 first what do you get?

OpenStudy (anonymous):

which is (pi)r^3=48

OpenStudy (anonymous):

Now since (pi) is irrational just show it as follows r^2 = 48/pi.....Now what would r=.

OpenStudy (anonymous):

sorry, it's r v-cubed, not squared. that stays the same tho right?

OpenStudy (anonymous):

*r cubed

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

okay. so r^3=15.27887454

OpenStudy (anonymous):

You are correct but I would prefer r = cube root of 64/(pi)

OpenStudy (anonymous):

cube root of r and 15.27887454

OpenStudy (anonymous):

so r = 2.481401964?

OpenStudy (anonymous):

I guess I do not have a calculator

OpenStudy (anonymous):

okay. thanks again. :) you've been a really big help today!

OpenStudy (anonymous):

HAve a good one..

OpenStudy (anonymous):

you too!